OpenStudy (lsdxmdmaddicthc):

Can someone help me solve this? Not looking for an answer, just an explanation! :) This is a Compound Interest Formula, I'm being asked to solve for the Interest Rate (represented by r). 2,147.39 = 1,852.1(1+(r/12))^60 I don't know how to isolate r. Any help is greatly appreciated It's Advanced Algebra with Financial Applications.

OpenStudy (lsdxmdmaddicthc):

\[2,147.39 = 1,852.1(1+\frac{ r }{ 12 })^{60}\]

OpenStudy (baru):

\[[(2147.39/1852.1)^\frac{ 1 }{ 60 } - 1] \times 12\] you will need a calculator

OpenStudy (lsdxmdmaddicthc):

Can you explain the steps involved? I have a calculator. I want to understand how to do that.

OpenStudy (lsdxmdmaddicthc):

I never learned how to deal with exponents in equations, and my teacher isn't providing assistance.

OpenStudy (baru):

you do not know how to get further than this, am i right \[2147.39/1852.1 = (1+ r/12)^6\]

OpenStudy (baru):

sorry the exponant is 60

OpenStudy (lsdxmdmaddicthc):

I can't eliminate the exponent.

OpenStudy (lsdxmdmaddicthc):

I don't know how to do the converse of the exponent to the other side of the equation.

OpenStudy (baru):

the trick is: you are allowed to "raise" both sides of the eqation to the same power

OpenStudy (lsdxmdmaddicthc):

?

OpenStudy (lsdxmdmaddicthc):

Ohhh I think I know what you mean Tell me if this is right? (2,147.39*1,852.1)^60 = (1+r/12)^60

OpenStudy (baru):

not quite: (2,147.39/1852.1)^(1/60) = ((1+r/12)^60)^(1/60) i'm sorry that must be quite hard to read

OpenStudy (baru):

i'll explain it differently

OpenStudy (baru):

do u know what it means when I say "square both sides of the equation?" (talking in genral, not this problem)

OpenStudy (lsdxmdmaddicthc):

Yes a^2 = a^2

OpenStudy (baru):

do you understand if I say "take the square root of both sides"

OpenStudy (lsdxmdmaddicthc):

Yes \[\sqrt{a} =\sqrt{a}\]

OpenStudy (baru):

now what you should undertand is : \[\sqrt{x+y} = (x+y)^{1/2} \] they mean the same thing

OpenStudy (lsdxmdmaddicthc):

Okay

OpenStudy (baru):

in this question (i'm simplifying) its something like \[x+y= (a+b)^{60}\]

OpenStudy (baru):

and i'm asking you to take the "sixtieth root on both sides"

OpenStudy (baru):

\[(x+y)^{1/60}=((a+b)^{60})^{1/60}\]

OpenStudy (lsdxmdmaddicthc):

x+y^(60/1) = (a+b)^60)^60/1

OpenStudy (lsdxmdmaddicthc):

Err yeah 1/60 Sorry typo

OpenStudy (baru):

not 60/1 its 1/60 square root: 1/2 sixtieth root: 1/60

OpenStudy (baru):

oh lol ok

OpenStudy (baru):

now use the rule \[((a)^{m})^n= (a)^{m \times n}\]

OpenStudy (lsdxmdmaddicthc):

So 1.1594 ^(1/60) = ((1+r/12)^60)^(1/60)

OpenStudy (baru):

yep

OpenStudy (lsdxmdmaddicthc):

THANK YOU SO MUCH!!! Are you a Senior in HS?

OpenStudy (baru):

nope...college

OpenStudy (lsdxmdmaddicthc):

Awesome! Best of luck to you :) Thanks again!

OpenStudy (baru):

sure :)

OpenStudy (baru):

you too